Circular No. 9076

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 S3 (LEMMON) An apparently asteroidal object discovered in the course of the Mt.

Lemmon survey (discovery observation tabulated below) has been found to be diffuse on follow-up CCD images taken by W. H. Ryan and E. V. Ryan on Sept. 26.2 UT with the Magdalena Ridge Observatory 2.4-m f/8.9 reflector, following posting on the Minor Planet Center’s ‘NEOCP’ webpage.

2009 UT  α_{2000}  δ_{2000}  Mag.
Sept.24.16207  18^{h}01^{m}23.^{s}44  +33^{o}14′06′′.3  20.9

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-S122.

T = 2011 Dec. 25.349 TT  ω = 129.^{o}389
Ω = 223.179 }2000.0
q = 6.68049 AU  i = 59.962
2000 WT_{169} AND 2003 YU_{179}

K. S. Noll, Space Telescope Science Institute (STScI); W. M. Grundy, Lowell Observatory; S. D. Benecchi, STScI; and H. A. Levison, Southwest Research Institute, report the detection of binary companions to the transneptunian objects 2000 WT_{169} (cf. MPECs 2000-X31, 2002-B11) and 2003 YU_{179} (cf. MPEC 2006-H29), from four dithered observations of each object that were made using the Hubble Space Telescopes Wide Field Planetary Camera 2 through the F606W filter starting on 2008 Sept. 24.3825 and 27.3823 UT, respectively. For 2000 WT_{169}, a second component was found at −0′′.069 ± 0′′.003 in α and −0′′.050 ± 0′′.001 in δ, which was fainter than the primary by 0.43 magnitude. The 2003 YU_{179} secondary was found at −0′′.039 ± 0′′.008 in α and −0′′.054 ± 0.005 in δ, being 1.3 magnitudes fainter than the primary.

COMET 225P/2009 Q3 (LINEAR) Corrigendum. On IAUC 9068, in the title, the designation should read

‘P/2009 Q3’.

2009 September 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green